• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name bogdanw

    Log entry time 16:07:47 on November18,2009

    Entry number 300799

    keyword=shift summary

    run is smooth but HRS-R Q2 is down
    8:10 beam up to 100 \mu A
    8:15 beam up to 104 \mu A
    8:55 beam OFF for high beam preparation
    9:00 lost HRS-l Q1 - helium level is down
    9:05 helium level back above 50%
    9:15 target on empty position, MCC send beam as large as 120
    9:40 MCC completed study, hall in control access
    run flipper - incremented slug number - become 5
    call MCC, they put IHWP out
    run flipper again without incrementing slug number
    Now log entry and status are consistent.
    12:25 tech completed work in the Hall
    HRS-L Q1 is ramping up
    13:00 HRS-L Q1 trips, Jack is going back to the Hall
    15:00 Jack is adusting Q1 cryo flow
    15:20 beak back, check on BeO is OK, harp scan is OK,
    raster 6x6 (MCC) and 2 muA beam ok holey carbon is OK.
    DAQ run for spot++ check is OK.